OpenStudy (anonymous):

What is the ordered pair of positive integers (a,b) for which a/b is a reduced fraction and x=[a(pi)]/b is the least positive solution of the equation: (2cos8x -1) (2cos4x -1) (2cos2x -1) (2cosx -1)=1?

OpenStudy (anonymous):

is that \((2\cos^8x-1)(2\cos^4x-1)(2\cos^2x-1)(2\cos x-1)\)?

OpenStudy (anonymous):

no \[(2\cos8x -1)(2\cos4x -1)(2\cos2x -1)(2cosx -1) =1\]

OpenStudy (anonymous):

0 can definitely be a value for x, except that 0 cannot be expressed as a fraction...

OpenStudy (anonymous):

this is damn tricky... I suppose it's \((1,1)\) by the fact that it yields \(x=\pi\) which is a trivial solution to the equation given \(\cos k\pi=0,k\in\mathbb{Z}\)

OpenStudy (anonymous):

2 pi works too

OpenStudy (anonymous):

cos 2pi = cos0

OpenStudy (anonymous):

oops what the hell did I just write

OpenStudy (anonymous):

I meant \(\cos k\pi=(-1)^k\) hehe

OpenStudy (anonymous):

no worries, i just saw the cosine graph and realize that

OpenStudy (anonymous):

since x=2pi, then (a,b) is (2,1)

OpenStudy (anonymous):

but then 2/1 isn't a reduced fraction... ><

OpenStudy (anonymous):

2/1 is a reduced fraction actually :-p

OpenStudy (anonymous):

oh.... but 2/1 should be simplified to 2

OpenStudy (anonymous):

Do you think that is correct? (a,b)--> (2,1)

OpenStudy (anonymous):

yes, indeed, and \(2/1\) is the number \(2\) in lowest \(p/q\) form. well let me think. I know that \(\cos(\pi n)=(-1)^n\) so let \(z=-1\) and we can write$$(2z^8-1)(2z^4-1)(2z^2-1)(2z-1)$$but hmm I wonder how to prove no \(\dfrac{a}b\pi\) works for \(a<b\)

OpenStudy (anonymous):

Thanks for looking at this with me!! Appreciate it~ =D

OpenStudy (anonymous):

I think I've got it

OpenStudy (anonymous):

so obvious in retrospect ;-)$$(2\cos(8x)-1)(2\cos(4x)-1)(2\cos(2x)-1)(2\cos(x)-1)=1$$so$$(2\cos(8x)-1)(2\cos(4x)-1)(2\cos(2x)-1)(4\cos^2(x)-1)=2\cos(x)+1$$recall that \(\cos^2(x)=\dfrac12(1+\cos(2x))\) ergo \(4\cos^2(x)=2+2\cos(2x)\) hence \(4\cos^2(x)-1=2\cos(2x)+1\) and we have:$$(2\cos(8x)-1)(2\cos(4x)-1)(2\cos(2x)-1)(2\cos(2x)+1)=2\cos(x)+1\\(2\cos(8x)-1)(2\cos(4x)-1)(4\cos^2(2x)-1)=2\cos(x)+1$$observe this effect continues to cascade:$$(2\cos(8x)-1)(2\cos(4x)-1)(2\cos(4x)+1)=2\cos(x)+1\\(2\cos(8x)-1)(4\cos^2(4x)-1)=2\cos(x)+1\\(2\cos(8x)-1)(2\cos(8x)+1)=2\cos(x)+1\\4\cos(8x)^2-1=2\cos(x)+1\\2\cos(16x)+1=2\cos(x)+1\\\cos(16x)=\cos(x)$$

OpenStudy (anonymous):

so we know that when \(\cos(16x)=\cos(x)\) it means that \(16x=2\pi k\pm x\) for integer \(k\). Observe:$$17x=2\pi k\\x=\frac{2\pi k}{17}$$or$$15x=2\pi k\\x=\frac{2\pi k}{15}$$in both cases it is obvious that we choose \(k=1\) and \(2/17<2/15\) ergo \((2,17)\) is our answer

OpenStudy (anonymous):

I am very expressed by your solution! Well done!!

OpenStudy (anonymous):

I was able to follow everything except I am a little confused on how did you get from step one to step two

OpenStudy (anonymous):

Multiplying by \(2\cos(x)+1\):$$(2\cos(x)-1)(2\cos(x)+1)=4\cos^2(x)-1$$

OpenStudy (anonymous):

that is I multiplied both sides by \(2\cos(x)+1\) and then simplified the LHS as above

OpenStudy (anonymous):

okay, how did you figure out that you have to multiply (2cosx +1) ?

OpenStudy (anonymous):

just wondering.....

OpenStudy (anonymous):

It was an initial guess of mine but I didn't inspect too closely... came back later and realized it worked :-p

OpenStudy (anonymous):

=D Wonderful~

OpenStudy (anonymous):

Well, thank you so much for figuring this one out with me!!! Appreciate it!!! =D

OpenStudy (anonymous):

NP -- glad I could help

OpenStudy (anonymous):

@oldrin.bataku Sorry for the trouble I wanted some clarification on the last 3 steps, as I was re-writing this on paper, I realize something different than your solution: \[4\cos^28x -1 = 2\cos x +1\] \[4(\cos^28x)-1=4(\frac{\cos 8x +1 }{ 2 }) =2\cos8x +1 \] then should it be \[2\cos 8x +1 =2\cos x +1\] then \[\cos 8x = \cos x\] Thus the solution will be 2/9 or 2/7 and (a,b) ---> (2,9)

OpenStudy (anonymous):

are you sure \(\cos^2(8x)=\dfrac12(\cos(8x)+1)\)? ;-)

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

I see my mistake now :P thanks again~

OpenStudy (anonymous):

mmm.... just one last justification, please, @oldrin.bataku I couldn't understand why 16x=2πk±x

OpenStudy (anonymous):

whenever you have \(\cos(x)=\cos(y)\) recall that we know \(\cos(x)=\cos(-x)\) so \(y\) could be \(\pm x\) and further we know that \(\cos(x+2\pi k)=\cos(x),k\in\mathbb{Z}\) i.e. that y could also be \(2\pi k\pm x\)

OpenStudy (anonymous):

essentially if we know two values \(x,y\) have the same cosine then we know they are (up to a difference that is an integer multiple of \(2\pi\) as \(\cos\cdot\) is periodic) equal up to sign change

OpenStudy (anonymous):

draw it: